What amount of substance containing 60% NaCl, 37% KCl should be weighed out for analysis so that after the action of 25ml of 0.1N AgNO3 solution, excess of Ag+ is back titrated with 5ml of NH4SCN solution? Given that 1ml of NH4SCN = 1.1ml of AgNO3.

Grade:11

1 Answers

Amisha Jaiswal
6 years ago
1ml of NH4SCN =1.1ml of agno35ml of NH4SCN=1.1×5=5.5mlAgNO3 LEFT- 25-5.5=19.5NO. OF MOLES(AGNO3) =0.1×19.5=n×1000n= 1.95/10^-3Now, 1 mole of nacl and 1 mole of kcl requires 1 mole of AGNO3NO.OF MOLES OF AGNO3 =NO. OF MOLES OF NACL=NO. OF MOLES OF KCLNACL> 0.00195=m/Mm= 0.114gMass of kcl = 0.14560+37%= 97%= 0.2592100%= 0.2592/97×100= 0.267g